THIRD-DEGREE PRICE DISCRIMINATION A seller faces two groups of buyers: Pi-16- Q1 and P2-24-Q2. Marginal cost is constant at $4 and fixed costs are zero a. Assuming that resale of the good by consumers is impossible, find profit-maximizing quantities and prices under 3rd-degree price discrimination. No need to calculate profit. Show graphs and math. Suppose someone argues Under this outcome Pi and P2 differ, so this cannot be profit maximizing since a seller could always transfer a unit from the low price market to the high price market and make a profit in doing so. What is wrong with this argument? BE BRIEF (ONE sentence) b. Suppose price discrimination is not possible, so all buyers face the same price. Find profit-maximizing total quantity, price, and profit. Note: make sure you show ALL work. Show this result on a graph. Your graph should include the quantity at which kink in the market demand occurs along with all relevant information about the MR curve (such as values at Qkink) c. Suppose that consumers can resell the good from one market to the next at a cost of S2 per unit. Using the Lagrangian method, find the profit maximizing quantities and prices. Show work.
Answer #1

a)  

P 24-01 f16- a. 2. Pray Mazini.zakion Pmoslom Max P aPa 02TC T HANE TR(24-8) G 2- . TR160,-6 hh: 16-20a、 2 MR2 2 4-2a2 162G- 司12。2G, ィー6 れ.SCo山 16 1よ1b 211

If seller transfers a unit from lower priced market to higher priced market, then demand equations changes and sum of MR OF 1 AND MR OF 2 are not equal to MC. So, not possible.

b) Following from (a), the seller cannot discriminate now. So, he must charge equal price in both markets.

Total demand (Q) + Q1 + Q2 = 16 - P + 24 - P = 40 - 2P

INVERSE DEMAND (P) = (40 - Q ) /2 = 20 - 0.5Q

MC = 4 (GIVEN)

TOTAL REVENUE (P*Q) = 20Q - 0.5 Q2 .  

So, MR = 20 - Q

At equlibrium, MC = MR

or, 4 = 20 - Q

or, Q* = 16 Now, P = 20 -0.5(16) = 16 - 8 = 8 or, P*=8

PROFIT = TR - TC = 20(16) - 0.5(162) - 4(16) = 320 - 128 - 64 = 128.

>MR 16 2

c) max R(q)

subject to pi = R(q) - C(q) >= pi* ---------------------------------------------------- (1)

Lagrangian function : L = R(q) + lambda [ R(q) - C(q) - pi*] ;  pi* is min acceptable profit.

Kuhn- tucker conditions are : (1) dL/dq = R'(q) + lambda [ R'(q) - C'(q) - pi ] <= 0

(2) q >= 0 q(dL/dq) = 0

(3)  lambda >= 0 lambda(dL/dlambda) = 0

(4) dl/dlambda = R(q) - C(q) - pi* >= 0  

Assuming pi* = 85.75 < profit of 128

Unrestricted Max R'(q) = 200 at Q* = 20 and pi = 120

Equality 1 is

20q - 0.5q2 - (4q+2q) = 85.75

or, 14q - 0.5q2 - 85.75 = 0

or, q2 -28q +171.5 = 0

Using sridhar acharya formula to solve the quadatic equation, we get

q = 9,19

therefore,p = 15.5, 10.5